Okay, I Promise This is the last math problem because my limit is 4.
I Understood the last one , but once I ran upon this one I Got confused again.
x + 4y =-19
-3x - 12y = 57
The answers are:dependent (-3,-4) inconsistent (-3,4)
When the problems threw in the dependent and inconsistent answers, I couldn't figure out the answer. Thank You So Much!

Answer by fractalier(2101)

Okay, cause I'm going home now...from
x + 4y = -19
-3x - 12y = 57
we can multiply the top equation by -3 and get
-3x - 12y = 57,
exactly the same as the other equation. Hence this system is dependent, they represent the same line and have an infinite number of solutions.